What is the exact circumference of a circle with a diameter of 13 inches?
26p inches
6.5p inches
13p inches
42.25p inches
169p inches
I was doing 13*3.14=40.84 but as you can see this is not one of my selections so I am not sure where I am going wrong.

Question #2 want to double check this one

Find the volume of the pyramid with a rectangular base shown below
h=11in
w=12in
l=10in

120 cubic inches
440 cubic inches
528 cubic inches
336 2/3 cubic inches
1320 cubic inches

11*12*10=1320 in




Answer by stanbon(75887)

What is the exact circumference of a circle with a diameter of 13 inches?
26p inches
6.5p inches
13p inches
42.25p inches
169p inches
I was doing 13*3.14=40.84 but as you can see this is not one of my selections so I am not sure where I am going wrong.
------------------------
Since pi is an endless decimal form your answer
is approximate. They want the exact answer which
is 13(pi).
They have listed that answer as 13p in you option list.
----
Correct answer: 13p
==============================================================

Question #2 want to double check this one
Find the volume of the pyramid with a rectangular base shown below
h=11in
w=12in
l=10in
---
V = (1/3)Bh
B = 12*10 = 120 in^2
h = 11
---
V = (1/3)120*11 in^3
V = 40*11
V = 440 in^3
===================
Cheers,
Stan H.
===================
120 cubic inches
440 cubic inches
528 cubic inches
336 2/3 cubic inches
1320 cubic inches
11*12*10=1320 in